# PDMS elastic properties: influence of fabrication protocol and test method

### v1

@@ Acknowledgements
H.V. acknowledges the support of the Department of Defense (DOD) through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. Funding agencies were not involved in study design; in the collection, analysis and interpretation of data; in the writing of the report; or in the decision to submit the article for publication.
Appendix A Connectedness of existing literature
Litmaps software was used to create a connectedness diagram of all the studies used in this review shown in Figure A1. [4] is seen to be particularly commonly cited with work on PDMS.
Figure A1: Included studies show citation links between articles. Markers are sized based on relevance within the network and sorted chronologically from left to right, and vertically on a log scale by total citations in any field per the Litmaps Database [64]
Appendix B Additional data
[…]
S55-10 50+5.6:1 2 3.65 $\pm$ 0.11 0.326 $\pm$ 0.01
S50-00a 50+0 :1 4 5.71 $\pm$ 0.16 0.319 $\pm$ 0.01
S50-00b 50+0 :1 3 4.71 $\pm$ 0.51 0.321 $\pm$ 0.01
S50-00c 50+0 :1 2 5.47 $\pm$ 0.29 0.342 $\pm$ 0.007
S50-10 45+5 :1 3 7.38 $\pm$ 0.37 0.366 $\pm$ 0.02
S50-20 40+10:1 3 7.24 $\pm$ 0.34 0.395 $\pm$ 0.006
S45-00a 45+0 :1 1 11.9 $\pm$ 0.24 0.448 $\pm$ 0.02
S45-00b 45+0 :1 4 10.6 $\pm$ 0.81 0.414 $\pm$ 0.03
S40-00 40+0 :1 4 14.4 $\pm$ 1.8 0.460 $\pm$ 0.02
Table 2: Additional data presented in this work with associated composition and fit parameters for shear modulus and initial defect size.
Appendix C Included studies
Author Test Type Cure Temp [C] Cure Time [hr:min] Mix Ratio [base+oil: curing agent]
Chockalingam 2021 [53] VCCE 100 2:00 variable, Si Oil
Raayai-Ardakani 2019 [17] VCCE 40 72:00 variable
Raayai-Ardakani 2019 [52] VCCE 40 72:00 variable
This work VCCE 100 2:00 variable, Si Oil
[…]
Milner 2021 Thesis [56] Cavitation Rheology 25 144:00 variable
Yang 2019 [57] Cavitation Rheology 25 144:00 variable, Si Oil
[…]
Upadhyay 2019 [58] Simple Shear 60 3:00 10:1
[…]
Fuard 2008 [8] Tension 100 1:30, 2:00 variable
Johnston 2014 [40] Tension† 25-200 variable 10:1
Khanafer 2008 [65] Tension 65 12:00 6-10:1
Kim 2011 [66] Tension 5-15:1
Liu 2009 [39] Tension† 100 or 200 variable 10:1
Mills 2008 [41] Tension‡ 150 12:00 10:1
Moučka 2021 [44] Tension† 25-150 variable 10:1, Si Oil
Schneider 2008 [36] Tension 150 0:15 10:1
Seghir 2015 [42] Tension 10 or 160 2:00, 144:00 variable
Upadhyay 2021 [38] Tension 60 3:00 10:1
Wang 2019 [21] Tension 65 4:00 10-30:1
[…]
Carrillo 2005 [46] Compression & Nanoindentation 25 336:00 10-30:1
[…]
Cheng 2011 [48] Nanoindentation 65 1:30 10:1
Gupta 2007 [50] Nanoindentation 25 0:20,2:40 variable
Liao 2010 [51] Nanoindentation 85 2:00 variable
Mata 2005 [45] Nanoindentation 95 0:30 5.7-21:1
Patel 2019 [19] Nanoindentation 70 5:00 10:1
Peng 2011 [20] Nanoindentation 60 20:00 50:1
[…]
Du 2010 [60] Beam Bending 65 1:30 10:1
[…]
Thangawng 2007 [61] Membrane 110 0:15 10:1
Table 3: Included studies. Tension tests performed per the ASTM D412 standard are indicated with †while tests per ASTM D 5045-99 are indicated with ‡.

### v2

@@ Acknowledgements
We acknowledge the partial support of our work through Office of Naval Research grant N000142312530, and the support from the National Science Foundation under award number CMMI-1942016.. H.V. acknowledges the support of the Department of Defense through the National Defense Science & Engineering Graduate Fellowship Program. Funding agencies were not involved in study design; in the collection, analysis and interpretation of data; in the writing of the report; or in the decision to submit the article for publication.
Notes and references
  • Ji et al. 2022 Y. Ji, A. Dagro, G. Dorgant, D. Starr and J. Wilkerson, Experimental Mechanics, 2022, 62, 799–812.
  • Malone et al. 2018 F. Malone, E. McCarthy, P. Delassus, P. Fahy, J. Kennedy, A. Fagan and L. Morris, Cardiovascular engineering and technology, 2018, 9, 489–502.
  • Varner et al. 2023 H. Varner, G. P. Sugerman, M. K. Rausch and T. Cohen, Journal of the Mechanical Behavior of Biomedical Materials, 2023, 143, 105901.
  • McDonald and Whitesides 2002 J. C. McDonald and G. M. Whitesides, Accounts of chemical research, 2002, 35, 491–499.
  • Xia and Whitesides 1998 Y. Xia and G. M. Whitesides, Annual review of materials science, 1998, 28, 153–184.
  • Raj M and Chakraborty 2020 K. Raj M and S. Chakraborty, Journal of Applied Polymer Science, 2020, 137, 48958.
  • Lötters et al. 1997 J. C. Lötters, W. Olthuis, P. H. Veltink and P. Bergveld, Journal of Micromechanics and Microengineering, 1997, 7, 145.
  • Wu et al. 2015 J. Wu, R. Wang, H. Yu, G. Li, K. Xu, N. C. Tien, R. C. Roberts and D. Li, Lab on a Chip, 2015, 15, 690–695.
  • Wu et al. 2010 W.-Y. Wu, X. Zhong, W. Wang, Q. Miao and J.-J. Zhu, Electrochemistry communications, 2010, 12, 1600–1604.
  • Zhang and Hutchens 2021 B. Zhang and S. B. Hutchens, Soft Matter, 2021, 17, 6728–6741.
  • Raayai-Ardakani and Cohen 2019 S. Raayai-Ardakani and T. Cohen, Extreme Mechanics Letters, 2019, 31, 100536.
  • Kothari et al. 2020 M. Kothari, Z. Lemon, C. Roth and T. Cohen, Soft Matter, 2020, 16, 9838–9843.
  • Patel et al. 2019 H. Patel, C. Yang, H. Lee and A. A. Pelegri, ASME International Mechanical Engineering Congress and Exposition, 2019, p. V009T11A031.
  • Peng et al. 2011 X. Peng, J. Huang, H. Deng, C. Xiong and J. Fang, Measurement Science and Technology, 2011, 22, 027003.
  • Wang et al. 2019 Z. Wang, C. Xiang, X. Yao, P. Le Floch, J. Mendez and Z. Suo, Proceedings of the National Academy of Sciences, 2019, 116, 5967–5972.
  • Jung et al. 2012 H.-C. Jung, J.-H. Moon, D.-H. Baek, J.-H. Lee, Y.-Y. Choi, J.-S. Hong and S.-H. Lee, IEEE Transactions on Biomedical Engineering, 2012, 59, 1472–1479.
  • Ariati et al. 2021 R. Ariati, F. Sales, A. Souza, R. A. Lima and J. Ribeiro, Polymers, 2021, 13, 4258.
  • Niu et al. 2007 X. Niu, S. Peng, L. Liu, W. Wen and P. Sheng, ADVANCED MATERIALS-DEERFIELD BEACH THEN WEINHEIM-, 2007, 19, 2682.
  • Chen et al. 2015 D. Chen, F. Chen, X. Hu, H. Zhang, X. Yin and Y. Zhou, Composites Science and Technology, 2015, 117, 307–314.
  • Kowol et al. 2022 P. Kowol, S. Bargmann, P. Görrn and J. Wilmers, Extreme Mechanics Letters, 2022, 54, 101724.
  • Jiao et al. 2021 Q. Jiao, M. Shi, T. Yin, Z. Suo and J. J. Vlassak, Extreme Mechanics Letters, 2021, 48, 101433.
  • Brown et al. 2005 X. Q. Brown, K. Ookawa and J. Y. Wong, Biomaterials, 2005, 26, 3123–3129.
  • Kim et al. 2007 J. Kim, J. Park, S. Yang, J. Baek, B. Kim, S. H. Lee, E.-S. Yoon, K. Chun and S. Park, Lab on a Chip, 2007, 7, 1504–1508.
  • Tanaka et al. 2007 Y. Tanaka, K. Sato, T. Shimizu, M. Yamato, T. Okano and T. Kitamori, Lab on a Chip, 2007, 7, 207–212.
  • Fuard et al. 2008 D. Fuard, T. Tzvetkova-Chevolleau, S. Decossas, P. Tracqui and P. Schiavone, Microelectronic Engineering, 2008, 85, 1289–1293.
  • Huh et al. 2010 D. Huh, B. D. Matthews, A. Mammoto, M. Montoya-Zavala, H. Y. Hsin and D. E. Ingber, Science, 2010, 328, 1662–1668.
  • Feinberg et al. 2007 A. W. Feinberg, A. Feigel, S. S. Shevkoplyas, S. Sheehy, G. M. Whitesides and K. K. Parker, Science, 2007, 317, 1366–1370.
  • Nawroth et al. 2012 J. C. Nawroth, H. Lee, A. W. Feinberg, C. M. Ripplinger, M. L. McCain, A. Grosberg, J. O. Dabiri and K. K. Parker, Nature Biotechnology, 2012, 30, 792–797.
  • Chen et al. 2018 J. Chen, J. Zheng, Q. Gao, J. Zhang, J. Zhang, O. M. Omisore, L. Wang and H. Li, Applied Sciences, 2018, 8, 345.
  • Wolf et al. 2018 M. P. Wolf, G. B. Salieb-Beugelaar and P. Hunziker, Progress in Polymer Science, 2018, 83, 97–134.
  • Zhou et al. 2010 J. Zhou, A. V. Ellis and N. H. Voelcker, Electrophoresis, 2010, 31, 2–16.
  • ABBASI et al. 2001 F. ABBASI, H. MIRZADEH and A.-A. KATBAB, Polymer international, 2001, 50, 1279–1287.
  • Teixeira et al. 2021 I. Teixeira, I. Castro, V. Carvalho, C. Rodrigues, A. Souza, R. Lima, S. Teixeira and J. Ribeiro, AIMS Materials Science, 2021, 8, .
  • Murphy et al. 2020 E. C. Murphy, J. H. Dumont, C. H. Park, G. Kestell, K.-S. Lee and A. Labouriau, Journal of Applied Polymer Science, 2020, 137, 48530.
  • Dow 2024 Dow, SYLGARD™ 184 Silicone Elastomer, 2024.
  • Heinrichs et al. 2018 V. Heinrichs, S. Dieluweit, J. Stellbrink, W. Pyckhout-Hintzen, N. Hersch, D. Richter and R. Merkel, PLoS One, 2018, 13, e0195180.
  • Palchesko et al. 2012 R. N. Palchesko, L. Zhang, Y. Sun and A. W. Feinberg, PloS one, 2012, 7, e51499.
  • Schweitzer et al. 2019 J. Schweitzer, S. Merad, G. Schrodj, F. Bally-Le Gall and L. Vonna, Journal of Chemical Education, 2019, 96, 1472–1478.
  • Kübler 1997 J. Kübler, Proceedings of the 21st Annual Conference on Composites, Advanced Ceramics, Materials, and Structures—B: Ceramic Engineering and Science Proceedings, 1997, pp. 155–162.
  • Quinn et al. 1994 G. D. Quinn, R. J. Gettings and J. J. Kübler, Proceedings of the 18th Annual Conference on Composites and Advanced Ceramic Materials—B: Ceramic Engineering and Science Proceedings, 1994, pp. 846–855.
  • Davies et al. 1992 P. Davies, H. Kausch, J. Williams, A. Kinloch, M. Charalambides, A. Pavan, D. Moore, R. Prediger, I. Robinson, N. Burgoyne et al., Composites Science and Technology, 1992, 43, 129–136.
  • Blackman et al. 2003 B. Blackman, A. Kinloch, M. Paraschi and W. Teo, International journal of adhesion and adhesives, 2003, 23, 293–305.
  • Van Mier et al. 1997 J. Van Mier, S. Shah, M. Arnaud, J. Balayssac, A. Bascoul, S. Choi, D. Dasenbrock, G. Ferrara, C. French, M. Gobbi et al., Materials and Structures, 1997, 30, 195–209.
  • Ruiz-Moreno et al. 2020 A. Ruiz-Moreno, P. Hähner, L. Kurpaska, J. Jagielski, P. Spätig, M. Trebala, S.-P. Hannula, S. Merino, G. de Diego, H. Namburi et al., Nanomaterials, 2020, 10, 130.
  • Liu et al. 2005 C. Liu, Y. Qiu and K. Lian, Microfluidics, BioMEMS, and Medical Microsystems III, 2005, pp. 179–185.
  • Schneider et al. 2008 F. Schneider, T. Fellner, J. Wilde and U. Wallrabe, Journal of Micromechanics and Microengineering, 2008, 18, 065008.
  • Armani et al. 1999 D. Armani, C. Liu and N. Aluru, Technical Digest. IEEE International MEMS 99 Conference. Twelfth IEEE International Conference on Micro Electro Mechanical Systems (Cat. No. 99CH36291), 1999, pp. 222–227.
  • Upadhyay et al. 2021 K. Upadhyay, D. Spearot and G. Subhash, International Journal of Impact Engineering, 2021, 156, 103949.
  • Khanafer et al. 2009 K. Khanafer, A. Duprey, M. Schlicht and R. Berguer, Biomedical microdevices, 2009, 11, 503–508.
  • Kim et al. 2011 T. K. Kim, J. K. Kim and O. C. Jeong, Microelectronic Engineering, 2011, 88, 1982–1985.
  • Liu et al. 2009 M. Liu, J. Sun and Q. Chen, Sensors and Actuators A: Physical, 2009, 151, 42–45.
  • Johnston et al. 2014 I. D. Johnston, D. K. McCluskey, C. K. Tan and M. C. Tracey, Journal of Micromechanics and Microengineering, 2014, 24, 035017.
  • Mills et al. 2008 K. Mills, X. Zhu, S. Takayama and M. Thouless, Journal of materials research, 2008, 23, 37–48.
  • Seghir and Arscott 2015 R. Seghir and S. Arscott, Sensors and Actuators A: Physical, 2015, 230, 33–39.
  • ASTM 2021 ASTM, Standard Test Methods for Vulcanized Rubber and Thermoplastic Elastomers—Tension, Astm standard, 2021.
  • Moučka et al. 2021 R. Moučka, M. Sedlačík, J. Osička and V. Pata, Scientific reports, 2021, 11, 19090.
  • Mata et al. 2005 A. Mata, A. J. Fleischman and S. Roy, Biomedical microdevices, 2005, 7, 281–293.
  • Carrillo et al. 2005 F. Carrillo, S. Gupta, M. Balooch, S. J. Marshall, G. W. Marshall, L. Pruitt and C. M. Puttlitz, Journal of materials research, 2005, 20, 2820–2830.
  • Wang et al. 2014 Z. Wang, A. A. Volinsky and N. D. Gallant, Journal of Applied Polymer Science, 2014, 131, .
  • Cheng 2011 C. Cheng, PhD thesis, The University of Texas at Dallas, 2011.
  • Cao et al. 2005 Y. Cao, D. Yang and W. Soboyejoy, Journal of Materials Research, 2005, 20, 2004–2011.
  • Gupta et al. 2007 S. Gupta, F. Carrillo, C. Li, L. Pruitt and C. Puttlitz, Materials Letters, 2007, 61, 448–451.
  • Liao et al. 2010 Q. Liao, J. Huang, T. Zhu, C. Xiong and J. Fang, Mechanics of Materials, 2010, 42, 1043–1047.
  • Raayai-Ardakani et al. 2019 S. Raayai-Ardakani, Z. Chen, D. R. Earl and T. Cohen, Soft matter, 2019, 15, 381–392.
  • Chockalingam et al. 2021 S. Chockalingam, C. Roth, T. Henzel and T. Cohen, Journal of the Mechanics and Physics of Solids, 2021, 146, 104172.
  • Li et al. 2024 J. Li, Z. Xie, H. Varner, S. Chockalingam and T. Cohen, In preparation, 2024.
  • Milner and Hutchens 2021 M. P. Milner and S. B. Hutchens, Mechanics of Materials, 2021, 154, 103741.
  • Milner 2020 M. P. Milner, PhD thesis, University of Illinois at Urbana-Champaign, 2020.
  • Yang et al. 2019 S. Yang, D. Bahk, J. Kim, A. Kataruka, A. C. Dunn and S. B. Hutchens, International Journal of Fracture, 2019, 219, 89–99.
  • Upadhyay et al. 2019 K. Upadhyay, A. Bhattacharyya, G. Subhash and D. Spearot, Experimental Mechanics, 2019, 59, 733–747.
  • Nunes 2011 L. Nunes, Materials Science and Engineering: A, 2011, 528, 1799–1804.
  • Du et al. 2010 P. Du, I.-K. Lin, H. Lu and X. Zhang, Journal of Micromechanics and Microengineering, 2010, 20, 095016.
  • Thangawng et al. 2007 A. L. Thangawng, R. S. Ruoff, M. A. Swartz and M. R. Glucksberg, Biomedical microdevices, 2007, 9, 587–595.
  • Spearman 1961 C. Spearman, Studies in individual differences: The search for intelligence, Appleton-Century-Crofts, 1961, pp. 45–58.
  • Everitt and Skrondal 2010 B. S. Everitt and A. Skrondal, The Cambridge dictionary of statistics, Cambridge University Press, 2010.
  • Akoglu 2018 H. Akoglu, Turkish journal of emergency medicine, 2018, 18, 91–93.
  • Esteves et al. 2009 A. Esteves, J. Brokken-Zijp, J. Laven, H. Huinink, N. Reuvers, M. Van and G. De With, Polymer, 2009, 50, 3955–3966.
  • Bardelli et al. 2022 T. Bardelli, C. Marano, F. Briatico Vangosa et al., Express Polymer Letters, 2022, 16, 924–932.